a body is projected at an angle 60 with horizontal plane with k.E 'k' when the velocity makes 30 with x axiz what is its k.E?


-if t1,t2 are the time of flights for two complimentary angles then range( in terms of g,t1,t2) is?

Grade:Upto college level

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear akash

let initial velocity is u

and v is the velocity when it make angel 30

ucos60=vcos30

v=u/√3

 

k=1/2 m u2

u=√(2k/m)

so v=√(2k/3m)

now K .E =1/2 m v2

             =k/3 


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin